OpenStudy (savannah_noelle):

Please explain how to do this please? The equation is f(x)=4x+10. Assign any number to x. Using complete sentences, explain whether f(g(x)) and g(f(x)) will result in the same number.

OpenStudy (savannah_noelle):

@jim_thompson5910 @satellite73 Please help?

OpenStudy (anonymous):

what is \(g(x)\)?

OpenStudy (anonymous):

in general \(f(g(x))\) is not equal to \(g(f(x))\) where you given a specific \(g\) to use?

OpenStudy (anonymous):

help me out here...

OpenStudy (savannah_noelle):

This is what I was given. I'm on number 6.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

this must be the dumbest problem of the day no matter, we can still solve if you have \[f(x)=4x+10\] now lets pick a \(g\) i pick a simple one \(g(x)=x^2+1\)

OpenStudy (anonymous):

then we pick a number i pick 2 so \[f(2)=4\times 2+10=14\] and therefore \[g(f(2))=g(14)=14^2+1=197\]

OpenStudy (anonymous):

now \[g(2)=2^2+1=5\] and so \[f(g(2))=f(5)=4\times 5+10=30\]

OpenStudy (anonymous):

and we see that \[f(g(2))=30, g(f(2))=197\] and so \[f(g(2))\neq g(f(2))\]

OpenStudy (savannah_noelle):

Thank you so much! This helped a lot. (:

OpenStudy (anonymous):

yw

OpenStudy (savannah_noelle):

@satellite73 for this step: hen we pick a number i pick 2 so f(2)=4×2+10=14 how did you get 14?

OpenStudy (ybarrap):

It was just an error, you can correct. The point is that f(g) is not equal to g(f) generally. You will find this when you correct in his example. Question is pretty funny though.

OpenStudy (anonymous):

lol i made a mistake, that is how

OpenStudy (savannah_noelle):

Thanks @ybarrap @satellite73 (: lol yeah, the question is pretty lame. I ended up calling my teacher and his response was "God that question is so stupid, it could have definitely been re-worded."